OpenStudy (anonymous):

(FILL IN BOXES) Seth built a wooden step stool out of two rectangular prisms and two cubes. Before fastening the components together, he stained each component to give it its final color. To completely stain the bottom rectangular prism, Seth had to cover [ ] square inches of wood. For one of the cubes to be completely stained, he had to cover [ ] square inches of wood. To completely stain the top rectangular prism, he had to cover[ ] square inches of wood. Once the stool was fastened together, he applied a coat of sealant to all exposed surfaces of the stool, including the bottom, to cover a total of [ ]square inches.

OpenStudy (anonymous):

I can help u

OpenStudy (anonymous):

@Directrix

OpenStudy (anonymous):

First part had to cover 712 square inches of wood Second Part had to cover 96 square inches of wood last part covering a total of 1176 square inches

OpenStudy (anonymous):

thank you but there are 4 boxes

Directrix (directrix):

Bottom Rectangular Prism Total Area TA = p*h + 2*B where p is the perimeter of the base and h is the height of the prism and B is the area of the bases of the prism. (2*22 + 2*8)*6 + 2*(22*8) = 712 cubic inches

Directrix (directrix):

@killaray23 I agree with you on that.

OpenStudy (anonymous):

the question went away :(

OpenStudy (anonymous):

https://media.studyisland.com/userfiles/ok8frame3.gif If X = 3 inches, Y = 11 inches, W = 5 inches, and Z = 5 inches, what is the area of the object? 45 square inches 35 square inches 24 square inches 66 square inches

OpenStudy (anonymous):

i wana give u both medals but i cant im gonna post this one on a different one:)

OpenStudy (anonymous):

ok txs

Directrix (directrix):

We are not finished. The phone rang and I had to stop working. Will return.

Directrix (directrix):

Total Area of One Cube (4*2 +4*2)*4 + 2* (4*4) = 96 cubic inches

Directrix (directrix):

Total Area of Top Prism (2*18 + 2*5)*2 + 2* (5*18) = 272 cubic inches

Directrix (directrix):

@killaray23 We differ here: To completely stain the top rectangular prism, he had to cover[ ] square inches of wood.

Directrix (directrix):

Note: Here is the same problem with the result for the top step missing. Interesting. http://openstudy.com/updates/53337f68e4b0c30c52c37e17

Directrix (directrix):

No work given there ^^^^

Directrix (directrix):

Fourth Part Once the stool was fastened together, he applied a coat of sealant to all exposed surfaces of the stool, including the bottom, to cover a total of [ ]square inches. Total surface area of separate prisms and subtract regions no longer exposed after assembly. (272 + 96 + 712) - 4*(4*4) - 4*(4*4) = 952 cubic inches

Directrix (directrix):

@killaray23 Need to ask about this: last part covering a total of 1176 square inches